OpenStudy (katielong):

algebra: solve x^2+3x (find 2 solutions)

OpenStudy (phi):

you need an equation not just an expression. An equation has an = sign in it. Do you mean \[ x^2 +3x = 0 \] ?

OpenStudy (katielong):

yes, sorry

OpenStudy (phi):

do you know how to "factor out" an x ? there is the "distributive law" x(x+3) can be written as x^2 +3x factor out the x to undo the "distribution" in other words, \[ x^2 +3x = 0 \\ x(x+3)= 0\]

OpenStudy (katielong):

yeah :)

OpenStudy (phi):

can you find the 2 solutions ?

OpenStudy (katielong):

not really.. x=0/(x+3) ? or x+3=0 so x=-3?

OpenStudy (phi):

yes. You can check your answer: try x=0 in x^2 + 3x if x=0, we have 0*0 + 3*0 = 0+0 = 0 it works try x=-3: -3*-3 + 3*-3= +9 - 9 = 0 it works Finding the answer uses this idea: if you have 2 numbers a*b=0 then you know that either a is 0 or b is 0 (or both are 0) you have x * (x+3) =0 so either x is 0 or (x+3) is 0 you then find either x=0 or x= -3 (and obviously x can't be both 0 and -3 at the same time)

OpenStudy (katielong):

okay, thanks alot, thats really clear!